# Learning Robust and Correct Controllers Guided by Feasibility-Aware Signal Temporal Logic via BarrierNet

### v2

@@ Abstract
Control Barrier Functions (CBFs) have emerged as a powerful tool for enforcing safety in optimization-based controllers, and their integration with Signal Temporal Logic (STL) has enabled the specification-driven synthesis of complex robotic behaviors. However, existing CBF–STL approaches typically rely on fixed hyperparameters and myopic, per–time step optimization, which can lead to overly conservative behavior, infeasibility near tight input limits, and difficulty satisfying long-horizon STL tasks. To address these limitations, we propose a feasibility-aware learning framework that embeds trainable, time-varying High Order Control Barrier Functions (HOCBFs) into a differentiable Quadratic Program (dQP). Our approach provides a systematic procedure for constructing time-varying HOCBF constraints for a broad fragment of STL and introduces a unified robustness measure that jointly captures STL satisfaction, QP feasibility, and control-bound compliance. Three neural networks—InitNet, RefNet, and an extended BarrierNet—collaborate to generate reference inputs and adapt constraint-related hyperparameters automatically over time and across initial conditions, reducing conservativeness while maximizing robustness. The resulting controller achieves STL satisfaction with strictly feasible dQPs and requires no manual tuning. Simulation results demonstrate that the proposed framework maintains high STL robustness under tight input bounds and significantly outperforms fixed-parameter and non-adaptive baselines in complex environments.
@@ I INTRODUCTION
Autonomous and robotic systems are often required to meet mission objectives that extend well beyond basic stability or invariance requirements. In practical scenarios such as surveillance, for instance, an unmanned aircraft may need to periodically collect information from a designated area, revisit a charging station for a minimum duration at regular intervals, and consistently steer clear of restricted airspace. To formally capture such time- and event-dependent requirements, temporal logics—most notably Linear Temporal Logic (LTL) [baier2008principles] and Signal Temporal Logic (STL) [maler2004monitoring]—have become standard specification languages due to their rich expressivity and clear semantics.
@@
In this work, we address the control of dynamical systems subject to requirements expressed in STL. STL provides a rich language for specifying time-dependent requirements over real-valued signals. It supports both Boolean semantics, which indicate whether a trajectory satisfies a formula, and quantitative (robustness) semantics [donze2010robust], which assign a real value reflecting the degree of satisfaction or violation. Prior research has shown that enforcing STL constraints can be formulated as an optimization problem in which the robustness metric appears either in the objective or as part of the constraints. These formulations have been tackled using Mixed Integer Linear Programming (MILP) [raman2014model, sadraddini2015robust] as well as gradient-based optimization methods [pant2017smooth, mehdipour2019arithmetic, gilpin2020smooth]. Although effective, these approaches require considerable computation, limiting their practicality for real-time or online control.
@@
Recent works have also employed Control Barrier Functions (CBFs) to enforce STL specifications on system trajectories [lindemann2018control, garg2019control, xiao2021high2, liu2024auxiliary]. CBFs have been extensively used in the controls community to guarantee safety by ensuring forward invariance of a prescribed safe set [ames2016control, ames2014control], and they can be combined with Control Lyapunov Functions (CLFs) to additionally encode stability objectives. In both cases, the CBF and CLF appear as inequality constraints on the control input, and the resulting controllers are typically synthesized through quadratic programs (QPs), which are computationally efficient and suitable for real-time implementation. Building on this foundation, the CBF framework has been extended to address higher relative-degree constraints [xiao2021high, nguyen2016exponential], mixed relative-degree constraints [liu2025auxiliary] and adaptive CBF formulations [xiao2021adaptive, liu2023auxiliary]. However, since the CBFs or hyperparameters involved in CBFs are typically selected manually, poor selection can lead to excessive conservativeness, and the resulting QPs may become infeasible due to conflicts between the CBF constraints and the input bounds. Although online adjustment of these hyperparameters can improve feasibility, it adds computational overhead and compromises practicality for real‐time applications [liu2025auxiliary, rabiee2025soft, parwana2025rate, breeden2023compositions].
@@
Learning-based methods can perform most of the computation offline during training, which enables real-time control during execution. In contrast to CBF–QP controllers, many STL-guided learning approaches do not formulate control synthesis as a constrained optimization problem either during training or at run time. For example, model-based Reinforcement Learning (RL) learns a policy by maximizing STL robustness [yaghoubi2019worst, leung2022semi], and Q-learning has also been applied to STL control synthesis in [aksaray2016q, venkataraman2020tractable]. In addition, the authors of [kalagarla2021model] formulate the STL control synthesis as a constrained Markov Decision Process (cMDP), which yields a computable lower bound on the STL satisfaction probability. More recently, the authors of [meng2025tgpo] propose Temporal Grounded Policy Optimization (TGPO), which learns STL-satisfying policies by decomposing the temporal specification into stage-wise rewards and training a hierarchical RL policy accordingly. Since these policies are trained solely through robustness-based objectives and are directly executed after training, they do not exhibit numerical infeasibility caused by conflicting constraints. Nevertheless, these learning-based approaches cannot guarantee that the resulting policy will satisfy the specification. Violations may occur for two reasons. First, neural network training may converge to a suboptimal local solution under complex dynamics or specifications, leading to undesired behavior. Second, even if a policy satisfies the specification on training data, it may still fail for unseen initial conditions or environments.
@@
To better ensure STL satisfaction, recent works integrate CBFs into learning-based controllers [liu2021recurrent, liu2023safe]. In both methods, however, CBF constraints are not enforced during policy optimization: [liu2021recurrent] trains an RNN to imitate trajectories that satisfy STL specifications generated offline under CBF constraints, while [liu2023safe] uses CBFs only to guarantee constraint-compliant data collection when learning a probabilistic model. In both cases, the learned policy itself does not encode STL correctness and must be corrected at execution via a runtime CBF–QP filter. This separation creates a mismatch between training and execution, making STL satisfaction dependent on the feasibility of the online QP, which may fail when CBF-based constraints conflict with system dynamics or input bounds. Conversely, directly enforcing CBF constraints during training often makes the optimization infeasible and prematurely halts learning. Thus, a key challenge is to enforce CBF constraints consistently during both training and execution while accounting for conflicts between constraint enforcement and system or input limits to preserve feasibility and maintain STL satisfaction.
@@
In this paper, we address STL satisfaction by enforcing CBF-based constraints during both training and execution. We adopt a model-based RL framework in which trajectory rollouts are generated by repeatedly solving a differentiable QP (dQP) whose last layer augments a neural network controller with CBF constraints, incorporating specification correctness directly into policy learning rather than handling it externally. We train three neural networks to generate time-varying hyperparameters in the dQP cost and constraints: (i) InitNet, which initializes hyperparameters based on the initial condition; (ii) RefNet, which updates cost-related hyperparameters along the rollout; and (iii) an extended BarrierNet architecture, inspired by [xiao2021barriernet], which adapts the HOCBF constraint hyperparameters. These networks are optimized via a unified robustness metric that jointly captures STL satisfaction and QP feasibility, improving performance while accounting for potential infeasibility. After training, the networks provide hyperparameters online for execution, enabling specification-compliant control under new initial conditions. By jointly optimizing specification robustness and feasibility, the resulting controller remains feasible even under tight input bounds or complex STL tasks. Overall, the method avoids manual CBF tuning, reduces over-conservativeness via hyperparameter learning, and enables feasible STL enforcement during both training and execution without requiring pre-generated reference inputs as in [xiao2021barriernet].
@@
This work substantially extends our prior conference paper [liu2023learning], which enforced CBF constraints in a dQP using BarrierNet to improve model-based RL for STL satisfaction. The previous formulation optimized only STL robustness and ignored dQP feasibility during training, often causing infeasible QP constraints under tight input limits or complex dynamics, and prematurely terminating learning. It also used CBF hyperparameters that remained constant within each rollout, limiting their ability to adapt over the time horizon and restricting robustness improvement. In this paper, we introduce a unified robustness measure that incorporates both STL satisfaction and QP feasibility, guiding learning toward solutions that remain feasible even in challenging scenarios. In addition, the CBF hyperparameters become time-varying along each rollout, enabling adaptive hyperparameter tuning and further reducing conservativeness while improving robustness. Simulation results further demonstrate that the proposed method achieves higher robustness under strict input limits and challenging STL specifications.
@@ II Preliminaries
where $\mathbf{x}(t)\in\mathbb{R}^{n}$ denotes the state and $\mathbf{u}(t)\in\mathcal{U}\subseteq\mathbb{R}^{q}$ denotes the control input. The functions $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ and $g:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n\times q}$ are assumed to be locally Lipschitz. The input set $\mathcal{U}$ is a compact hyper‐rectangle, i.e., $\mathbf{u}_{min}\leq\mathbf{u}\leq\mathbf{u}_{max}$, where the inequality is interpreted element‐wise. Without loss of generality, we set the initial time to 0. The initial condition $\mathbf{x}(0)=\mathbf{x}_{0}$ is sampled from a set $\mathcal{X}_{0}\subseteq\mathbb{R}^{n}$ according to a probability density function $P:\mathcal{X}_{0}\rightarrow\mathbb{R}$. We consider system trajectories over a compact time horizon $[0,T]$. A signal $\mathbf{x}:[0,T]\rightarrow\mathbb{R}^{n}$ is a solution of the system (1) if it is absolutely continuous and satisfies the dynamics for all $t\in[0,T]$. A partial trajectory over $[0,t]$ is denoted as $\mathbf{x}_{0:t}$.
@@
where the controller depends on state histories and can be implemented using a recurrent architecture such as a Recurrent Neural Network (RNN) [goodfellow2016deep].
@@ II-A Signal Temporal Logic (STL)
Signal Temporal Logic (STL) [maler2004monitoring] is defined over real-valued signals $\mathbf{x}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}^{n}$, such as trajectories generated by (1). In this work, we focus on a fragment of STL characterized by the following syntax:
$\phi$ $\coloneqq\top\ |\ \mu\ |\ \neg\mu\ |\ \phi_{1}\land\phi_{2}\ |\ \phi_{1}\vee\phi_{2},$ (4a)
@@
where $\phi$ and $\varphi$ denote STL formulae; $\phi_{1}$ and $\phi_{2}$ are formulae of class $\phi$, and $\varphi_{1}$ and $\varphi_{2}$ are formulae of class $\varphi$; $\top$ denotes the logical true; $\mu$ represents an atomic predicate of the form $h(\mathbf{x})\geq 0$, where $h:\mathbb{R}^{n}\rightarrow\mathbb{R}$; $\neg$, $\land$, and $\vee$ denote Boolean negation, conjunction, and disjunction, respectively; $F$ and $G$ are the temporal operators eventually and always, respectively; and $t_{a},t_{b}]$ is a time interval such that $t_{a}<t_{b}$.
We write $(\mathbf{x},t)\models\varphi$ to indicate that the signal $\mathbf{x}$ satisfies the formula $\varphi$ at time $t$. A formal definition of the qualitative semantics of STL is provided in [maler2004monitoring]. Informally, $F_{[t_{a},t_{b}]}\phi$ holds if $\phi$ becomes true at some time in $t_{a},t_{b}]$,” whereas $G_{[t_{a},t_{b}]}\phi$ holds if $\phi$ is true at all times in $t_{a},t_{b}]$.” Boolean operators are interpreted in the standard way. Compared with the full STL syntax in [maler2004monitoring], the fragment in (4b) excludes the temporal until operator, nested temporal operators such as “eventually always,” and Boolean disjunction $\vee$ in order to retain a tractable structure for control and monitoring. Despite this restriction, the fragment remains expressive enough to capture a broad class of practically relevant temporal requirements, including safety and reachability specifications with explicit timing constraints.
STL is also equipped with quantitative semantics, known as robustness, which assigns a real value that measures the degree to which a signal satisfies a formula $\varphi$. Several robustness metrics have been proposed in the literature [donze2010robust, mehdipour2019arithmetic, gilpin2020smooth, varnai2020robustness]. In this work, we adopt the smooth robustness introduced in [liu2023robust], which is differentiable almost everywhere and can be readily embedded in learning-based algorithms. This robustness measure is sound, in the sense that its value is positive if and only if the STL formula is satisfied. We denote the robustness of $\varphi$ at time $t$ with respect to a signal $\mathbf{x}$ by $\rho(\varphi,\mathbf{x},t)$. We also define the time horizon of an STL formula $\varphi$, denoted by $hrz(\varphi)$, as the earliest future time required to determine its satisfaction and robustness. Throughout this paper, we evaluate trajectories of (1) only over the time horizon of the specification, i.e., we set $T=hrz(\varphi)$.
@@ II-B Time-Varying High Order Control Barrier Function
In this subsection, we introduce time-varying High Order Control Barrier Functions (HOCBFs) [xiao2021high2]. We begin by recalling the definition of a class $\kappa$ function:
Definition 1 (Class $\kappa$ function [Khalil:1173048]).
@@
A set $\mathcal{C}(t)\subset\mathbb{R}^{n}$ is forward invariant for system (1) if its solutions for some $\mathbf{u}\in\mathcal{U}$ starting from any $\boldsymbol{x}(0)\in\mathcal{C}(0)$ satisfy $\boldsymbol{x}(t)\in\mathcal{C}(t),\forall t\geq 0.$
Definition 4 (Time-varying High Order Control Barrier Function (HOCBF) [xiao2021high]).
Let $\psi_{1}(\mathbf{x},t),\ldots,\psi_{m}(\mathbf{x},t)$ be defined by (5) and $\mathcal{C}_{1}(t),\ldots,C_{m}(t)$ be defined by (6). A differentiable function $b(\mathbf{x},t)$ is a High Order Control Barrier Function (HOCBF) with relative degree $m$ with respect to system (1) if there exist differentiable class $\kappa$ functions $\alpha_{i}$, $i=1,\ldots,m$, such that
@@
for all $(\mathbf{x},t)\in\mathcal{C}_{1}(t)\cap\mathcal{C}_{2}(t)\cap\ldots\cap\mathcal{C}_{m}(t)\times[0,T]$. In (7), $L_{f}^{m}$ denote $m$-th order Lie derivatives along $f$; $L_{g}$ is the Lie derivative along $g$. We use $O(\cdot)=\sum_{i=1}^{m-1}L_{f}^{i}(\alpha_{m-1}\circ\psi_{m-i-1})(\mathbf{x},t)$ and assume that $L_{g}L_{f}^{m-1}b(\mathbf{x},t)\mathbf{u}\neq 0$ on the boundary of set $\mathcal{C}_{1}(t)\cap...\cap\mathcal{C}_{m}(t)$.
Theorem 1 (Safety Guarantee [xiao2021high]).

### v3

# Feasibility-aware Learning of Robust Temporal Logic Controllers using BarrierNet

@@ Abstract
Control Barrier Functions (CBFs) have been used to enforce safety and task specifications expressed in Signal Temporal Logic (STL). However, existing CBF-STL approaches typically rely on fixed hyperparameters and per-step optimization, which can lead to overly conservative behavior, infeasibility near tight input limits, and difficulty satisfying long-horizon STL tasks. To address these limitations, we propose a feasibility-aware learning framework that constructs trainable, time-varying High Order Control Barrier Function (HOCBF) constraints and hyperparameters that guarantee satisfaction of a given STL specification. We introduce a unified robustness measure that jointly captures STL satisfaction, constraint feasibility, and control-bound compliance, and propose a neural network architecture to generate control inputs that maximize this robustness. The resulting controller guarantees STL satisfaction with strictly feasible HOCBF constraints and requires no manual tuning. Simulation results demonstrate that the proposed framework maintains high STL robustness under tight input bounds and significantly outperforms fixed-parameter and non-adaptive baselines in complex environments.
@@ I INTRODUCTION
Autonomous and robotic systems are often required to meet mission objectives that extend well beyond basic stability or invariance requirements. In practical scenarios such as surveillance, for instance, an unmanned aircraft may need to periodically collect information from a designated area, revisit a charging station for a minimum duration at regular intervals, and consistently steer clear of restricted airspace. To formally capture such time- and event-dependent requirements, temporal logics—most notably Linear Temporal Logic (LTL) [5] and Signal Temporal Logic (STL) [25]—have become standard specification languages due to their rich expressivity and easily interpretable semantics.
@@
In this work, we address the control of dynamical systems subject to requirements expressed in STL, which provides a rich language for specifying time-dependent requirements over real-valued signals. STL supports both Boolean semantics, which indicate whether a trajectory satisfies a formula, and quantitative (robustness) semantics [7], which assign a real value reflecting the degree of satisfaction or violation. Prior research has shown that enforcing STL constraints can be formulated as an optimization problem in which the robustness metric appears either in the objective or in the constraints. These formulations have been tackled using Mixed Integer Linear Programming (MILP) [32, 33] as well as gradient-based optimization methods [29, 26, 9]. Although effective, these approaches require considerable computation, limiting their practicality for real-time or online control.
@@
Recent works have also employed Control Barrier Functions (CBFs) to enforce STL specifications on system trajectories [16, 8, 36, 18]. CBFs have been extensively used in the controls community to guarantee safety by ensuring forward invariance of a prescribed safe set [3, 2], and they can be combined with Control Lyapunov Functions (CLFs) to additionally encode stability objectives. In both cases, the CBF and CLF appear as inequality constraints on the control input, and the resulting controllers are typically synthesized through quadratic programs (QPs), which are computationally efficient and suitable for real-time implementation. Building on this foundation, the CBF framework has been extended to address higher relative-degree constraints [38, 28], mixed relative-degree constraints [19] and adaptive control [37, 17]. However, since the CBFs or hyperparameters involved in CBFs are typically selected manually, poor selection can lead to excessive conservativeness, and the resulting QPs may become infeasible due to conflicts between the CBF constraints and the input bounds. Although online adjustment of these hyperparameters can improve feasibility, it adds computational overhead and compromises practicality for real‐time applications [19, 31, 30, 6].
@@
Learning-based methods can perform most of the computation offline during training, which enables real-time control during execution. In contrast to CBF–QP controllers, many STL-guided learning approaches do not formulate control synthesis as a constrained optimization problem either during training or at run time. For example, model-based Reinforcement Learning (RL) learns a policy by maximizing STL robustness [40, 15], and Q-learning has also been applied to STL control synthesis in [1, 35]. In addition, the authors of [11] formulate the STL control synthesis as a constrained Markov Decision Process (cMDP), which yields a computable lower bound on the STL satisfaction probability. More recently, the authors of [27] propose Temporal Grounded Policy Optimization (TGPO), which learns STL-satisfying policies by decomposing the temporal specification into stage-wise rewards and training a hierarchical RL policy accordingly. Since these policies are trained solely through robustness-based objectives and are directly executed after training, they do not exhibit numerical infeasibility caused by conflicting constraints. Nevertheless, these learning-based approaches cannot guarantee that the resulting policy will satisfy the specification. Violations may occur for two reasons. First, neural network training may converge to a suboptimal local solution under complex dynamics or specifications, leading to undesired behavior. Second, even if a policy satisfies the specification on training data, it may still fail for unseen initial conditions or environments.
@@
To better ensure STL satisfaction, recent works integrate CBFs into learning-based controllers [22, 23]. In both these works, however, CBF constraints can only guarantee the “globally” subformulas of an STL specification, and they are not enforced during policy optimization: [22] trains an RNN to imitate trajectories that satisfy STL specifications generated offline under CBF constraints, while [23] uses CBFs only to guarantee constraint-compliant data collection when learning a probabilistic model. In both cases, the learned policy itself does not encode STL correctness and must be corrected at execution via a runtime CBF–QP filter. This separation creates a mismatch between training and execution, making STL satisfaction dependent on the feasibility of the online QP, which may fail when CBF-based constraints conflict with system dynamics or input bounds. Conversely, directly enforcing CBF constraints during training often makes the optimization infeasible and prematurely halts learning. Thus, a key challenge is to enforce CBF constraints consistently during both training and execution while accounting for conflicts between constraint enforcement and system or input limits to preserve feasibility and maintain STL satisfaction.
@@
In this paper, we guarantee STL satisfaction by enforcing CBF-based constraints during both training and execution. We provide a general, algorithmic procedure to generate these CBFs given an STL formula. We adopt a model-based RL framework in which trajectory rollouts are generated by repeatedly solving a differentiable QP (dQP) whose last layer augments a neural network controller with CBF constraints, incorporating specification correctness directly into policy learning rather than handling it externally. We train three neural networks to generate time-varying hyperparameters in the dQP cost and constraints: (i) InitNet, which initializes hyperparameters based on the initial condition; (ii) RefNet, which updates cost-related hyperparameters along the rollout; and (iii) an extended BarrierNet architecture, inspired by [39], which adapts the HOCBF constraint hyperparameters. These networks are optimized via a unified robustness metric that jointly captures STL satisfaction and QP feasibility, improving performance while accounting for potential infeasibility. After training, the networks provide hyperparameters online for execution, enabling specification-compliant control under new initial conditions. By jointly optimizing specification robustness and feasibility, the resulting controller remains feasible even under tight input bounds or complex STL tasks. Overall, the method avoids manual CBF tuning, reduces over-conservativeness via hyperparameter learning, and enables feasible STL enforcement during both training and execution without requiring pre-generated reference inputs as in [39].
@@
This work substantially extends our prior conference paper [24], which enforced CBF constraints in a dQP using BarrierNet to improve model-based RL for STL satisfaction. The previous formulation optimized only STL robustness and ignored dQP feasibility during training, often causing infeasible QP constraints under tight input limits or complex dynamics, and prematurely terminating learning. It also used CBF hyperparameters that remained constant within each rollout, limiting their ability to adapt over the time horizon and restricting robustness improvement. In this paper, we introduce a unified robustness measure that incorporates both STL satisfaction and QP feasibility, guiding learning toward solutions that remain feasible even in challenging scenarios. In addition, the CBF hyperparameters become time-varying along each rollout, enabling adaptive hyperparameter tuning and further reducing conservativeness while improving robustness. Simulation results further demonstrate that the proposed method achieves higher robustness under strict input limits and challenging STL specifications.
@@ II Preliminaries
where $\mathbf{x}(t)\in\mathbb{R}^{n}$ denotes the state and $\mathbf{u}(t)\in\mathcal{U}\subseteq\mathbb{R}^{q}$ denotes the control input. The functions $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ and $g:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n\times q}$ are assumed to be locally Lipschitz. The input set $\mathcal{U}$ is a compact hyper‐rectangle, i.e., $\mathbf{u}_{min}\leq\mathbf{u}\leq\mathbf{u}_{max}$, where the inequality is interpreted element‐wise. Without loss of generality, we set the initial time to 0. The initial condition $\mathbf{x}(0)=\mathbf{x}_{0}$ is sampled from a set $\mathcal{X}_{0}\subseteq\mathbb{R}^{n}$ according to a probability density function $P:\mathcal{X}_{0}\rightarrow\mathbb{R}$. We consider system trajectories over a compact time horizon $[0,T]$. A signal $\mathbf{x}:[0,T]\rightarrow\mathbb{R}^{n}$ is a solution of system (1) if it is absolutely continuous and satisfies the dynamics for all $t\in[0,T]$. A partial trajectory over $[0,t]$ is denoted as $\mathbf{x}_{0:t}$.
@@
where the controller depends on state histories and can be implemented using a recurrent architecture such as a Recurrent Neural Network (RNN) [10].
@@ II-A Signal Temporal Logic (STL)
Signal Temporal Logic (STL) [25] is defined over real-valued signals $\mathbf{x}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}^{n}$, such as trajectories generated by (1). In this work, we focus on a fragment of STL characterized by the following syntax:
$\phi$ $\coloneqq\top\ |\ \mu\ |\ \neg\mu\ |\ \phi_{1}\land\phi_{2},$ (4a)
@@
where $\phi$ and $\varphi$ denote STL formulae; $\phi_{1}$ and $\phi_{2}$ are formulae of class $\phi$, and $\varphi_{1}$ and $\varphi_{2}$ are formulae of class $\varphi$; $\top$ denotes the logical true; $\mu$ represents an atomic predicate of the form $h(\mathbf{x})\geq 0$, where $h:\mathbb{R}^{n}\rightarrow\mathbb{R}$; $\neg$ and $\land$ denote Boolean negation and conjunction, respectively; $F$ and $G$ are the temporal operators eventually and always, respectively; and $t_{a},t_{b}]$ is a time interval such that $t_{a}<t_{b}$.
We write $(\mathbf{x},t)\models\varphi$ to indicate that the signal $\mathbf{x}$ satisfies the formula $\varphi$ at time $t$. A formal definition of the qualitative semantics of STL is provided in [25]. Informally, $F_{[t_{a},t_{b}]}\phi$ holds if $\phi$ becomes true at some time in $t_{a},t_{b}]$”, whereas $G_{[t_{a},t_{b}]}\phi$ holds if $\phi$ is true at all times in $t_{a},t_{b}]$”. Boolean operators are interpreted in the standard way. Compared with the full STL syntax in [25], the fragment in (4) excludes the temporal until operator, nested temporal operators such as “eventually always”, and Boolean disjunction $\vee$ in order to be able to convert the STL formula into CBFs. Despite this restriction, the fragment remains expressive enough to capture a broad class of practically relevant temporal requirements, including safety and reachability specifications with explicit timing constraints.
STL is also equipped with quantitative semantics, known as robustness, which assigns a real value that measures the degree to which a signal satisfies a formula $\varphi$. Several robustness metrics have been proposed in the literature [7, 26, 9, 34]. In this work, we adopt the smooth robustness introduced in [21], which is differentiable almost everywhere and can be readily embedded in learning-based algorithms. This robustness measure is sound, in the sense that its value is positive if and only if the STL formula is satisfied. We denote the robustness of $\varphi$ at time $t$ with respect to a signal $\mathbf{x}$ by $\rho(\varphi,\mathbf{x},t)$. We also define the time horizon of an STL formula $\varphi$, denoted by $hrz(\varphi)$, as the earliest future time required to determine its satisfaction and robustness. Throughout this paper, we evaluate trajectories of (1) only over the time horizon of the specification, i.e., we set $T=hrz(\varphi)$.
@@ II-B Time-Varying High Order Control Barrier Function
In this subsection, we introduce time-varying High Order Control Barrier Functions (HOCBFs) [36]. We begin by recalling the definition of a class $\kappa$ function:
Definition 1 (Class $\kappa$ function [12]).
@@
In practice, the class $\kappa$ functions $\alpha_{i}(\cdot)$ are parameterized by a set of scalar hyperparameters $p_{i}$ that regulate how aggressively the associated functions are enforced, e.g., through a simple scaling of the form $p_{i}\alpha_{i}(\cdot)$.
@@
A set $\mathcal{C}(t)\subset\mathbb{R}^{n}$ is forward invariant for system (1) if its solutions for some $\mathbf{u}\in\mathcal{U}$ starting from any $\mathbf{x}(0)\in\mathcal{C}(0)$ satisfy $\mathbf{x}(t)\in\mathcal{C}(t),\forall t\geq 0.$
Definition 4 (Time-varying High Order Control Barrier Function [38]).
Let $\psi_{1}(\mathbf{x},t),\ldots,\psi_{m}(\mathbf{x},t)$ be defined by (5) and $\mathcal{C}_{1}(t),\ldots,C_{m}(t)$ be defined by (6). A differentiable function $b(\mathbf{x},t)$ is a High Order Control Barrier Function (HOCBF) with relative degree $m$ with respect to (1) if there exists differentiable class $\kappa$ functions $\alpha_{i}$, $i=1,\ldots,m$, such that
@@
for all $(\mathbf{x},t)\in\mathcal{C}_{1}(t)\cap\mathcal{C}_{2}(t)\cap\ldots\cap\mathcal{C}_{m}(t)\times[0,T]$. In (7), $L_{f}^{m}$ denote $m$-th order Lie derivatives along $f$; $L_{g}$ is the Lie derivative along $g$. Here $O(\cdot)$ collects the remaining Lie derivatives along $f$ and partial derivatives with respect to $t$ of degree less than $m$, as well as mixed derivatives of total order less than or equal to $m$. Note that the mixed derivative terms vanish when $b(\mathbf{x},t)$ is additively separable, i.e., when $\frac{\partial^{2}b(\mathbf{x},t)}{\partial x\partial t}=0$, a condition satisfied by the CBFs considered in this paper. Here, We assume that $L_{g}L_{f}^{m-1}b(\mathbf{x},t)\mathbf{u}\neq 0$ on the boundary of set $\mathcal{C}_{1}(t)\cap...\cap\mathcal{C}_{m}(t)$.
Theorem 1 (Safety Guarantee [38]).